# Euclid Quick Data Release (Q1)

### v1

@@ 6.1 Inpainting of AGN
For our inpainting, we use a 55 pixel mask centred on the brightest pixel, following the constraints described in Sect. 4.2. Comparing the expected pixel values with the generated outputs, Fig. 9 shows the model’s performance. The assumption that the model will predict a fainter core than the original does appear to be true, with a near global bias for underpredicting as shown by the dashed line. The majority of sources behave in a consistent way with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original. This distinct cluster could provide sufficient candidates for AGN.
The horizontal bar in the leftmost panel of Fig. 9, where the maximum output value appears to equal one, is a side effect of Eq. 11. The values themselves are not exactly one, with 1% of the data fall within of having a maximum output of one. There does exist a convergence point in Eq. 11 , where when the maximum and minimum pixel values of an input image equal one, the output of the rescaling always equals one; however, such an image would have all its pixels valued one, and so is not present in the data. Although there is not a specific convergence point that explains this trend, the majority of the images that fall on this line occur when the network outputs a maximum pixel value close to one. However, there are many more images that have a network output close to one that are not close to this line, as well as images close to the line without a network output close to one. We have verified that both point-like and extended source appear within this line, therefore due to the low number of sources impacted, we leave more in-depth exploration to future work.
To further validate the output of the model and show it is producing realistic outputs, Fig. 19 shows how the ROC metric compares between input and output images. The significant overlap, especially with respect to peak of each distribution, shows that the model is accurately recreating the dynamics of the data.
@@
There is a clear distribution difference between the respective class candidates. Stars perform the most consistently across magnitudes, producing the highest MSEs. There is an inherent trend for how high the error of the predictions are with the relative brightness of the classes, with stars and the brightest QSOs featuring the highest errors across all magnitudes. This is then followed by QSOs at higher magnitudes, and the lowest magnitude AGN candidates. Finally, the galaxy sample, along with the faintest AGN candidates produce much lower errors. Since these classes are candidates, it could also highlight the presence of contamination within the AGN labels. However, the large number of DESI-detected galaxies showing vastly reduced errors implies that the model does develop a bias for recreating galaxies with little or no AGN component.
The second important finding is the trend between VIS magnitude and error rate across all classes. The brighter a source is, the higher its recorded MSE. This matches the initial assumptions of the method, where, as the sources increase in magnitude, differences in brightness values across galaxies become harder to differentiate. This leads to fainter potential AGN candidates looking more similar to a normal galaxy, causing the inpainting to reconstruct more accurately and therefore less likely to be included in our selection. However, the relationship between brightness and error is once again more influenced by the scale-variant nature of MSE. However, the lack of significant overlap for many of the selections with the galaxy class still make this metric useful.
To showcase how the performance of the inpainting method impacts the overall image, Fig. 11 presents the input, output and residuals for a selection of sources, including a close up of the masked area both before and after inpainting. It is clear to see how bright central cores can be entirely removed with inpainting, being replaced by regions that are consistent with the surrounding pixels. The model is also able to recreate images that do not feature any significant core accurately. This shows the model has a good internal representation of galaxies and how they should behave.
@@ 6.2 Max pixel difference ratio
Although reconstruction error allows stars and a subset of AGN to stand out from the rest, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes. Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output. As an example, if the reconstructed image ended up being the correct values but just shifted by one pixel, this would be penalised heavily under MSE. From the perspective of this paper, such a scenario would imply that the model knew a bright pixel was valid and therefore, should not be penalised. Another example would be if only the brightest pixel is different in the reconstruction. MSE would significantly reduce the impact of this measurement, potentially scoring similar to a source that gave small errors in every pixel within the mask.
@@
To ensure such cases are not overlooked, we measure the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction. As shown in Fig. 12, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores. The majority of the galaxy subset produce a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
@@
Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values. Many of the high density peaks remain consistent, even as the magnitude of the sources increase. This implies that this metric is more robust at detecting fainter AGN than MSE.
@@
The distribution also highlights the scale of the extreme values with many sources showing pixel values approaching 100 times greater than their counterparts. Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities. Sources producing the highest values for the ratio are those where the reconstruction produces a cutout that is significantly fainter than the original image. As can be seen in Fig. 20, this space is dominated by point-like objects. Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority, if not the entirety of the source. When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels. In these cases, the best match for empty space being more empty space.
@@
The behaviour of the previous examples match with the expected behaviour of the model and the initial assumptions of the paper. Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of the original pixels. However, there exists a subset of sources that exhibit the opposite behaviour and generate a much brighter collection of pixels. Figure 21 shows examples of such sources. When analysing the sources that exhibit this behaviour we find that they are overwhelmingly low S/N images. Given the difference in behaviour of these images during training, as shown in Fig. 7, it is clear this has had an impact on the model’s behaviour with these types of images.
Because we would like to limit this behaviour, we analyse how many of the unseen sources produce this type of output. The number of sources whose max pixel difference ratio is less than one for each of the four models discussed in Sect. 4, are shown in Table 2. The normalised loss helps significantly in reducing the number of sources exhibiting this behaviour. The normalised loss trained on the full data set was selected as our applied model due to it further reducing the number of outlier predictions.
@@ 6.3 Morphology-based performance
Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model is able to handle the complexities of different galaxy morphologies. Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al. (2025), we provide examples of edge-on (disk-edge-on_yes_fraction ), spiral (has-spiral-arms_yes_fraction ), and mergers (merging_merger_fraction ) in Figs. 22, 23, LABEL:, and 24, respectively.
These images show how our diffusion model and the inpainting pipeline effectively adapt to different morphologies. For edge-on in particular, the region of bright pixels within the masked area is often narrower than the mask itself. Despite this, the inpainting correctly preserves the orientation of the features rather than uniformly filling the entire masked area with bright pixels.
@@ 6.4 Creating a classifier
Due to the difference in distributions between the two metrics, any non-overlapping cases can be found when used in combination. As seen in Fig. 13, there is a cluster of sources in the top right. The key aspect of this cluster is its clear boundary from the DESI galaxy selection. This is vital for ensuring high reliability in classifications. We create a simple decision boundary between the two clusters of points, as shown in Fig. 13. Although a more fine-tuned boundary would certainly be possible, especially when incorporating a further machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics. The equation of our linear boundary is .
@@ 6.5 Variance of reconstructions
We perform the inpainting pipeline seven separate times on the same subset of the data. To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before. The mean and variance over a normalised by brightness MSE are shown in Fig. 14.
@@
The top figure shows that irrespective of magnitude, the model achieves similar relative median error across all images, highlighting the impact of the normalised loss throughout training. However, the difference in the scheduler over images of different quality and brightness is made evident in the bottom figure. The consistency of generated outputs vary drastically as fainter (and therefore lower S/N) images produce widely different pixels when resampled.
6.6 Comparing to other criteria
It is important to compare our AGN and QSO predictions to more traditional methods that are also able to utilise only the image data. Given the assumption of our model that sources containing an AGN will feature a bright central component with a sharp falloff in the galaxy’s light profile, comparing to Srsic flux and point spread function (PSF) flux measurements is crucial because they capture the key concepts of our assumptions. The Srsic flux helps quantify galaxy light distribution and differentiate an AGN’s central brightness from the galaxy’s structural brightness. PSF flux isolates the galaxy’s core, capturing point-like behaviour. The distribution of sources with these parameters is shown in Fig. 15. The equation of this linear boundary is .
@@ 6.6 Comparing to other criteria
Our second comparison is with the original rate of change metric used in Sect. 3.3. As we were able to see a clear distinction between the galaxy class and the others, it is important to verify how such a metric performs as a classifier. This is especially true given such a metric can be generated from the raw image, with no machine-learning model required. The boundary for this classifier is .
@@
The test set scores for each respective selection is shown in Fig. 16, with the collated results for stars, galaxies, AGN and QSOs shown in Fig. 25. All predictions are on images that were not used during training, nor used in the creation of any of the decision boundaries for the classifiers.
Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination, therefore our analysis focuses primarily on achieving sufficiently high recall. High recall ensures that our method is capable of identifying the broad range of AGN and QSO sources that are detectable through existing techniques. A high recall metric indicates that our classifier effectively captures the variety of sources each respective selection criterion is designed to detect.
While precision is also a consideration, our primary metric of success in this context is the diffusion model’s recall. Having the ability to recover a large fraction of known AGN and QSO types, across various selection methods, validates the robustness and comprehensiveness of our classifier. Due to differing constraints for each of the selections, we split the analysis to allow for a more representative comparison.
@@
The predictions from each classifier for whether a source is a normal galaxy is simply the inverse of the boundary’s original predictions. As Fig. 16 shows, all classifiers capture nearly all galaxies across every magnitude bin. The precision scores also match extremely closely between classifiers, indicating similar contaminants proving difficult, especially for the brighter sources. Such contaminants can be seen in Figs. 13 and 15, where stars and DESI-selected QSO fall outside the boundaries.
The ability to separate normal galaxies from other extragalactic or stellar objects has benefits for both groups of data. Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understand of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al. 2024).
@@
Due to the assumptions made for this method, the incidental capture of stars within our classifications is an expected outcome. The near-perfect recall shows that our classifier is able to capture stars well. This performance is beneficial when attempting to capture a high purity galaxy sample as discussed in the previous section. However, as other works such as Euclid Collaboration: Matamoro Zatarain et al. (2025) have shown, there are metrics to mitigate the contamination of stars. This is shown in practise by the high precision, highlighting how we are able to isolate this star selection from other point-like sources.
However, a significant number of stars are likely to remain in our sample. Therefore, for all subsequent selections, we apply the cut phz_star_prob_phz_class, to predictions of all classifiers to minimise the contamination of stellar sources.
@@
The Euclid photometric-based selections for QSOs include _ and _, as described in MZ25, and the Bisigello24-A 2-colour selection (Euclid Collaboration: Bisigello et al. 2024). This subset consider only sources with MUMAX_MINUS_MAG , a parameter that indicates point-likeness (Euclid Collaboration: Romelli et al. 2025).
@@
As these sources are assumed point-like, applying the above MUMAX_MINUS_MAG limit would be reasonable. However, due to the improvements in spatial resolution from DESI, at about , to Euclid’s now 02, some sources that appeared point-like through DESI, now appear extended. Therefore, to ensure all potential sources are found, we do not apply the restriction to this subset.
This selection provides the first evidence of significant differences within the candidates for each classifier. The inpainting classifier is able to capture a significant amount of the selection compared to the other classifiers. Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall. Whereas the diffusion model has higher coverage over the brightest sources (lowest magnitudes), the ROC classifer matched more candidates in the fainter sources.
@@
For the Gaia-detected QSOs, we use the Quaia selection (Storey-Fisher et al. 2024). The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25). We once again apply the MUMAX_MINUS_MAG restriction to the sources.
After the initial magnitude bin, our classifier was able to match with a high proportion of the Quaia-selected QSOs. The other two classifiers achieved equal recall scores, but in the faintest detected sources our classifier was able to match with an additional 20% of sources.
The particularly poor precision for this selection across all the classifiers is due to the considerable imbalance between positive and negative predictions. With only 0.5% of the Gaia-detected sources being Quaia candidates, selecting more than just the candidates is likely, significantly impacting the precision.
@@
Utilising WISE-detected sources allows us to verify our ability to select candidates that may have been obscured by dust. Using the selection criteria derived by Assef et al. (2018), we compare our candidates to the 90% reliability (R90) selection which is a high purity AGN sample.
@@
The main difference with the results of the X-ray-detected AGN is how the precision outperforms the recall across nearly every magnitude bin. This is due to low number of X-ray sources compared to the other selections. The recall performance between the diffusion model and the ROC classifier is very similar, with the PSF and Srsic flux classifier again struggling to match as many candidates as the other two.
Each of the selection criteria we compare with captures distinct subsets of the AGN and QSO populations, influenced by specific observational biases and methodological constraints. Therefore, it is only when collating these selections and by identifying trends and overlaps in the labels, that we can be confident that we sufficiently cover the spectrum of possible AGN.
Table 3 summarises the scores shown in Fig. 16. We combine the results into two magnitude bins 1820 and 2022.5, separating into brighter and fainter sources respectively. Due to similar scores in many of the selections, we highlight in bold any scores that outperform the other classifiers by atleast 0.05.

### v2

@@ 5.1 Inpainting of AGN
For our inpainting, we use a 55 pixel mask centred on the brightest pixel, following the constraints described in Sect. 3.1. Comparing the expected pixel values with the generated outputs, Fig. 5 shows the model’s performance. The assumption that the model will predict a fainter core than the original does appear to be true, with a near-global bias for underpredicting, as shown by the dashed line. The majority of sources behave in a consistent way, with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original. This distinct cluster could provide sufficient candidates for AGN.
The horizontal bar in the leftmost panel of Fig. 5, where the maximum output value appears to equal one, is a side effect of Eq. 11. The values themselves are not exactly one, with 1% of the data falling within of having a maximum output of one. There does exist a convergence point in Eq. 11 , where when the maximum and minimum pixel values of an input image equal one, the output of the rescaling always equals one; however, such an image would have all its pixels valued one, and so is not present in the data. Although no specific convergence point explains this trend, the majority of the images that fall on this line occur when the network outputs a maximum pixel value close to one. However, many more images have a network output close to one that is not close to this line, as well as images close to the line without a network output close to one. We have verified that both point-like and extended sources appear within this line; therefore, due to the low number of sources impacted, we leave more in-depth exploration to future work.
To further validate the model’s output and show it is producing realistic outputs, Fig. 19 shows how the ROC metric compares between input and output images. The significant overlap, especially with respect to the peak of each distribution, shows that the model is accurately recreating the dynamics of the data.
@@
There is a clear distribution difference between the respective class candidates. Brighter sources, particularly stars and luminous QSOs, produce higher errors, followed by fainter QSOs and the lowest-magnitude AGN candidates. Galaxies and the faintest AGN candidates show significantly lower errors, showing the model’s tendancy to reconstruct sources without strong AGN features more effectively. This indicates that the model has learned the anticipated bias introduced by the uneven distribution within the data set. While MSE is scale-dependant and influenced by the source’s VIS magnitude, the separation between the errors from the galaxy class and those of the other classes suggests it remains useful as a selection metric. The presence of low-error AGN candidates may also indicate contamination within the AGN labels.
It is acknowledged that because the model is trained on the full, unlabeled galaxy population, its prior distribution does contain information about AGN, even if they are rare. This could, in principle, lead the model to occasionally reconstruct AGN-like features, resulting in false negatives. However, the strong performance in separating known AGN and QSO populations, as shown in Figure 6, suggests that the prior is sufficiently dominated by non-active galaxies for this outlier detection framework to be effective.
To showcase how the performance of the inpainting method impacts the overall image, Fig. 7 presents the input, output and residuals for a selection of sources, including a close-up of the masked area both before and after inpainting. It is clear how bright central cores can be entirely removed with inpainting and replaced by regions consistent with the surrounding pixels. The model can also accurately recreate images that do not feature any significant core. This shows that the model has a good internal representation of galaxies and how they should behave.
@@ 5.2 Max pixel difference ratio
Although reconstruction error allows stars and a subset of AGN to stand out, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes. Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output. For example, if the reconstructed image ended up being the correct values but shifted by one pixel, this would be penalised heavily under MSE. From the perspective of this paper, such a scenario would represent a desirable outcome, as the model would correctly predict the presence of a bright, unresolved component, even if its exact location were shifted by a few pixels within the mask. Another example would be if only the brightest pixel is different in the reconstruction. MSE would significantly reduce the impact of this measurement, potentially scoring similarly to a source that gave minor errors in every pixel within the mask.
@@
To ensure such cases are not overlooked, we measure the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction. As shown in Fig. 8, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores. The majority of the galaxy subset produces a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
@@
Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values. Many of the high-density peaks remain consistent, even as the magnitude of the sources increases. This implies that this metric is more robust at detecting fainter AGN than MSE.
@@
The distribution also highlights the scale of the extreme values, with many sources showing pixel values approaching 100 times greater than their counterparts. Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities. Sources producing the highest values for the ratio are those where the reconstruction produces a cutout significantly fainter than the original image. As seen in Fig. 20, this space is dominated by point-like objects. Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority of the source, if not the entirety. When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels. In these cases, the point-like source contributed very little to the surrounding pixels, leading to the generation of only background noise without any source.
@@
The behaviour of the previous examples matches the model’s expected behaviour and the paper’s initial assumptions. Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of, the original pixels. However, a subset of sources exists that exhibit the opposite behaviour and generate a much brighter collection of pixels. Figure 21 shows examples of such sources. When analysing the sources that exhibit this behaviour, we found that they have overwhelmingly low S/N images. Further exploration of the impact of S/N on the diffusion process is shown in Sect. 6.1.
@@
The number of sources whose max pixel difference ratio is less than one for models trained on the two loss functions discussed in Sect. 2.3 are shown in Table 2. The normalised loss significantly reduces the number of sources exhibiting this behaviour. The normalised loss trained on the complete data set was selected as our applied model because it further reduced the number of erroneous predictions.
@@ 5.3 Morphology-based performance
Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model can handle the complexities of different galaxy morphologies. Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al. (2025), we provide examples of edge-on (disk-edge-on_yes_fraction ), spiral (has-spiral-arms_yes_fraction ), and mergers (merging_merger_fraction ) in Figs. 22, 23, LABEL:, and 24, respectively.
These images show how our diffusion model and the inpainting pipeline can effectively adapt to different morphologies. For edge-on in particular, the region of bright pixels within the masked area is often narrower than the mask itself. Despite this, the inpainting correctly preserves the orientation of the features rather than uniformly filling the entire masked area with bright pixels.
@@ 5.4 Creating a classifier
Due to the difference in distributions between the two metrics, any non-overlapping cases can be found when used in combination. As seen in Fig. 9, there is a cluster of sources in the top right. The key aspect of this cluster is its clear boundary from the DESI galaxy selection. This is vital for ensuring high reliability in classifications. We create a simple decision boundary between the two clusters of points, as shown in Fig. 9. Although a more fine-tuned boundary would be possible, especially when incorporating a secondary machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics. The equation of our linear boundary is .
@@ 5.5 Variance of reconstructions
7 samples are generated as part of a single application of the pipeline. To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before. The mean and variance over a normalised-by-brightness MSE are shown in Fig. 10.
@@
The top figure shows that irrespective of magnitude, the model achieves similar relative median error across all images, highlighting the impact of the normalised loss throughout training. However, the difference in the scheduler over images of different quality and brightness is evident in the bottom figure. The consistency of generated outputs varies drastically as fainter (and therefore typically lower S/N) images produce widely different pixels when resampled.
5.6 Comparing to other classifiers
It is important to compare our AGN and QSO predictions to more traditional methods that are also able to utilise only the image data.
Srsic and PSF Flux Measurements
To benchmark our non-parametric inpainting method against a more traditional approach, we construct a classifier based on parametric modelling of galaxy light profiles. We utilise morphological parameters sourced directly from the Euclid MER Q1 catalogue (Euclid Collaboration: Romelli et al. 2025). Specifically, we use the provided flux_vis_sersic_euclid and flux_vis_psf_euclid measurements. These parameters are derived by the MER pipeline using the SourceExtractor++ software (Bertin et al. 2020; Kümmel et al. 2020). For each detected source, the pipeline performs two separate and independent model fits. The flux_vis_psf_euclid value represents the total flux obtained by fitting a model of the instrumental PSF to the source image. The flux_vis_sersic_euclid value represents the total flux from fitting a two-dimensional Srsic profile, which is designed to model the extended light distribution of a galaxy. The underlying assumption for a classifier based on these parameters is that sources with a dominant, unresolved central component (such as a QSO or a Type-1 AGN) will be well-described by a PSF model, resulting in a high flux_vis_psf_euclid value. Conversely, normal resolved galaxies will be better described by the extended Srsic profile, leading to a high flux_vis_sersic_euclid value. By comparing the relative strength of these two flux measurements (as shown in Fig. 15), we can construct a simple linear classifier to distinguish between point-like and extended sources. This baseline method is computationally inexpensive as it relies on pre-computed catalogue values, but it is fundamentally limited to this simple morphological distinction and may struggle with identifying AGN in complex host galaxies or at lower relative luminosities.
The distribution of sources with these parameters is shown in Fig. 11. The equation of this linear boundary is .
ROC Metric
@@ 5.6 Comparing to other classifiers
Our second comparison is with the original rate of change metric used in Sect. 4.3. As a clear distinction between the galaxy class and the others is visible, it is important to verify how such a metric performs as a classifier. This is especially true given that such a metric can be generated from the raw image, and no machine-learning model or precomputed values are required. The boundary for this classifier is .
Performance against other candidates
Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination. Therefore our analysis focuses primarily on achieving sufficiently high recall. For an initial candidate selection method, such as the one proposed here, the primary goal is often to compile a large, inclusive sample for potential follow-up observations or more detailed multi-wavelength analysis. In this context, high recall is typically prioritised over high precision, which measures the purity of the selected sample. The scientific cost of missing a potentially interesting source (a false negative) is often considered higher than the operational cost of including a contaminant (a false positive) that can be filtered out in subsequent analysis stages. Therefore, our primary goal is to demonstrate the capability of our model to achieve high recall across a wide variety of AGN and QSO candidates which have been selected via different, more conventional methods and physical mechanisms.
Precision and Recall are defined as follows
For a given classification task (e.g., identifying DESI QSO objects):
TP (true positives) is the number of objects that belong to the class and are correctly predicted as belonging to it by the model.
FP (false positives) is the number of objects that do not belong to the class but are incorrectly predicted as belonging to it by the model.
FN (false negatives) is the number of objects that belong to the class but are not predicted as belonging to it by the model.
@@
The test set scores for each respective selection are shown in Fig. 12, with the collated results for stars, galaxies, AGN and QSOs shown in Fig. 25. All predictions are made from images that were not used during training nor used in the creation of any of the decision boundaries for the classifiers. Due to differing constraints for each selection, we split the analysis to allow for a more representative comparison.
@@
The predictions from each classifier for whether a source is a normal galaxy are simply the inverse of the boundary’s original predictions. As Fig. 12 shows, all classifiers capture nearly all galaxies across every magnitude bin. This behaviour is expected, as the selection criteria used by each classifier are explicitly designed to favour AGN-like features, thereby excluding typical galaxies by construction. The precision scores also match extremely closely between classifiers, indicating similar contaminants proving difficult, especially for the brighter sources. Such contaminants can be seen in Figs. 9 and 11, where stars and DESI-selected QSO fall outside the boundaries.
The ability to separate normal galaxies from other extragalactic or stellar objects benefits both groups of data. Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understanding of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al. 2025).
@@
Due to the assumptions made for this method, the incidental capture of stars within our classifications is an expected outcome. The near-perfect recall shows that our classifier is able to capture stars well. This performance is beneficial when attempting to capture a high-purity galaxy sample, as discussed in the previous section. However, as other works such as Euclid Collaboration: Matamoro Zatarain et al. (2025) have shown, there are metrics to mitigate the contamination of stars. As a significant number of stars are likely to remain in our sample, for all subsequent selections, we apply the cut phz_star_prob_phz_class, to predictions of all classifiers to minimise the contamination of stellar sources.
@@
The Euclid photometric-based selections for QSOs include _ and _, as described in MZ25, and the Bisigello24-A 2-colour selection (Euclid Collaboration: Bisigello et al. 2024). This subset considers only sources with MUMAX_MINUS_MAG , a parameter that indicates point-likeness (Euclid Collaboration: Romelli et al. 2025). As these selections have yet to go through a follow-up validation, they remain candidates and as such should not be seen as complete ground truth.
@@
As these sources are assumed point-like, applying the above MUMAX_MINUS_MAG limit would be reasonable. However, due to the improvements in spatial resolution from DESI, at about , to Euclid’s now 02, some sources that appeared point-like through DESI now appear extended. Therefore, to ensure all potential sources are found, we do not apply the restriction to this subset.
This selection provides the first evidence of significant differences in the suggested candidates for each classifier. The inpainting classifier is able to capture a significantly higher selection of lower magnitude sources compared to the other classifiers. Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall. The ROC classifier matched more candidates in the fainter sources.
@@
For the Gaia-detected QSOs, we use the Quaia selection (Storey-Fisher et al. 2024). The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25). We again apply the MUMAX_MINUS_MAG restriction to the sources.
After the initial magnitude bin, our classifier matched with a high proportion of the Quaia-selected QSOs. The other two classifiers achieved equal recall scores, but in the faintest detected sources, our classifier was able to match with an additional 20% of sources.
The particularly poor precision for this selection across all the classifiers is due to the considerable imbalance between positive and negative predictions. With only 0.5% of the Gaia-detected sources being Quaia candidates, selecting more than just the candidates is very likely, significantly impacting the precision.
@@
Utilising WISE-detected sources allows us to verify our ability to select candidates that may have been obscured by dust. Using the selection criteria derived by Assef et al. (2018), we compare our candidates to the 90% reliability (R90) selection, a high purity AGN sample.
@@
The main difference with the results of the X-ray-detected AGN and the other selections is that the precision outperforms the recall across nearly every magnitude bin. This is due to the low number of X-ray sources available. The recall performance between the diffusion model and the ROC classifier is very similar, with the PSF and Srsic flux classifier again struggling to match as many candidates as the other two.
Each of the selection criteria we compare captures distinct subsets of the AGN and QSO populations, influenced by specific observational biases and methodological constraints. Therefore, it is only when collating these selections and identifying trends and overlaps in the labels that we can be confident that we sufficiently cover the spectrum of possible AGN.
Table 3 summarises the scores shown in Fig. 12. We combine the results into two magnitude bins 18–20 and 20–22.5, separating into brighter and fainter sources, respectively. Due to similar scores in many of the selections, we highlight in bold any scores that outperform the other classifiers by at least 0.05.